# On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation

### v1

@@ 5 Conclusions
On this foundation, we develop analytical models for inventory parameters that represent the average behavior of the supply chain over an infinite time horizon, under parametric replenishment policies. Based on these models, we introduce optimization problems relevant to two key stakeholders: the constellation operator and the OOS provider. To capture the OOS provider’s decision-making context, we further introduce a functional relationship between OOS cost and responsiveness, characterized by a strictly decreasing curve with a knee point and two asymptotes parallel to the coordinate axes, implying cost and performance bounds. Finally, we formulate each stakeholder’s problem as an individual optimization model, and then integrate them into a bi-objective optimization framework to identify the family of Pareto-optimal (ideal) solutions. Finally, the developed models and optimization framework are applied to realistic large-scale constellation maintenance scenarios under various OOS-related parameters.
There are several promising avenues for future research. First, this study constrains the cost-responsiveness relationship to a specific functional form characterized by two shift and two shape parameters. Although this form can be applied to many scenarios, the resulting insights are not universal. Moreover, since these four parameters are defined based on current estimates, and OOS technology is still emerging, the relationship between cost and performance remains uncertain and likely to evolve. Therefore, further investigation into this relationship would enhance the relevance and generality of the results presented here.
Second, we assume that a recovered system returns to an “as-good-as-new” state. In practice, however, recovered satellites are likely to experience some degree of degradation. This factor could affect the effectiveness of OOS as part of the overall maintenance strategy, making it an important topic for future study.
Third, future research may also explore more detailed and realistic decision-making contexts and relevant solution concepts. Although the current solution concept, Pareto optimality, provides valuable theoretical insights, such idealized solutions are not always directly attainable in practice. Alternative frameworks may better represent real-world interactions. For instance, in situations where the constellation operator and OOS providers act in self-interest, or where multiple OOS providers compete for service opportunities, game-theoretic approaches such as the Nash equilibrium could yield more practical and policy-relevant insights.

### v2

@@ 6 Conclusions
On this foundation, we develop analytical models for inventory parameters that represent the average behavior of the supply chain over an infinite time horizon, under parametric replenishment policies. Based on these models, we introduce optimization problems relevant to two key stakeholders: the constellation operator and the OOS provider. To capture the OOS provider’s decision-making context, we further introduce a functional relationship between OOS cost and responsiveness, characterized by a strictly decreasing curve with a knee point and two asymptotes parallel to the coordinate axes, implying cost and performance bounds. We then formulate the interaction between the two stakeholders—the constellation operator and the OOS provider—as a bi-objective optimization problem to identify the family of Pareto-optimal solutions. The resulting models and optimization framework are applied to realistic large-scale constellation maintenance scenarios under various OOS-related parameter settings.
The case study yields three main findings. First, within a given scenario, the Pareto-optimal set is generally generated by an almost invariant maintenance strategy on the constellation operator’s side, while most variation along the Pareto frontier is driven by the provider-side pricing decision. Second, integrating OOS improves the maintenance system without degrading the required service levels, indicating that its value lies not only in the economic exchange between stakeholders but also in reducing dependence on conventional replacement through new satellite production and launch. Third, among the OOS-related parameters, the fraction of serviceable failure plays the most important structural role, as it is the only parameter that materially changes the constellation operator’s preferred maintenance architecture, while the remaining parameters mainly shift or reshape the Pareto frontier without substantially altering the operator-side decisions.
Future work should therefore proceed in three main directions. First, the cost-responsiveness relationship should be refined using more realistic, technology- and mission-specific models, especially as OOS-related technologies, policies, and market conditions continue to evolve. Second, the recovery process could be extended to account for imperfect restoration and state-dependent degradation. Third, the present bi-objective framework should be expanded to incorporate more realistic strategic interactions, such as bargaining and competition among multiple OOS providers.